OpenStudy (anonymous):

Please help :/ Find f^-1(x) if f(x)=5x-2/4 f^-1(x)= ?

OpenStudy (anonymous):

If f(x) = x + 4, find f(3). f(3) = 3 + 4 = 7 If g(x) = 2(x − 4), find the value of x if g(x) = 20. 2(x − 4) = 20 x - 4 = 10 x = 14

OpenStudy (anonymous):

So, a. (4x-5)/2 b. (4x+2)/5 c. (4x-2)/5 or d. (4x+5)/2?